# Characteristic Learning for Provable One Step Generation

### v3

@@ C.1. Auxiliary Properties
$=\frac{1}{(2\pi)^{d}\sigma\alpha_{t}}\frac{1}{\rho_{t}(x)}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big{)}\exp\Big{(}-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big{)}\,\mathrm{d}\nu(u)$
$=\frac{1}{(2\pi)^{d/2}}\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\int_{{\mathbb{R}}^{d}}\exp\Big{(}-\frac{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\sigma^{2}\alpha_{t}^{2}}\Big{\|}x_{1}-\frac{\sigma^{2}\beta_{t}x+\alpha_{t}^{2}u}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{\|}_{2}^{2}\Big{)}\,\mathrm{d}\nu(u),$
which can be viewed as a convolution of $\nu$ and a Gaussian distribution. Therefore, we can construct a random variable which obeys the conditional density $\rho_{1|t}(x_{1}|x)$ as
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
where $U$ and $\epsilon$ are two independent random variables sample from $\nu$ and $N(0,I_{d})$, respectively. This completes the proof. ∎
Lemma C.6 (Conditional expectation).
@@
Proof of Lemma C.6.
According to Lemma C.5, it holds that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
@@
Lemma C.7 (Conditional covariance).
@@
Proof of Lemma C.7.
It is straightforward from Lemma C.5 that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
@@
$=\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}{\mathbb{E}}[U]{\mathbb{E}}[U]^{T}+\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]x^{T}+\Big{(}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}xx^{T}.$
On the other hand, using Lemma C.5 deduces
${\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x]$ $=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]x^{T}+\Big{(}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}xx^{T}$
$\quad+\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}{\mathbb{E}}[UU^{T}]+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
@@
$=\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{)}^{2}\operatorname{Cov}(U)+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
According to Lemma C.5, the random variable $\|U\|_{\infty}\leq 1$ and thus $\operatorname{Cov}(U)\preceq dI_{d}$. Consequently,
@@
By a same argument as Lemmas C.6 and C.7, we also have the following inequality. See Gao et al. [2024, Lemma A.8] for a detailed proof.
Lemma C.8.
@@
C.2. Proof of Results in Section 3.2
@@
where the second equality holds from the definition of stochastic interpolant (2.1), and the last equality is due to Lemma C.2. This completes the proof. ∎
@@
$=\beta_{t}\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\Big{(}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x\Big{)}+\frac{\dot{\alpha}_{t}}{\alpha_{t}}x$
$=\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x$
where the second equality holds from the definition of stochastic interpolant (2.1), and the fourth equality follows from Lemma C.6. This completes the proof. ∎
@@
where the second inequality follows from Lemma C.6, and $M$ is a constant only depending on $d$ and $\sigma$. Applying Gronwall’s inequality [Evans, 2010, Section B.2] yields
@@
It follows from the definition of the probability flow $g^{*}$ (2.4) that
$\partial_{t}g^{*}(t,s,x)=-b^{*}(t,x)\quad\text{and}\quad\partial_{s}g^{*}(t,s,x)=b^{*}(s,x).$
Then we obtain the desired result according to Proposition 3.2. ∎
@@
(C.8) $\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}\preceq\nabla b^{*}(t,x)\preceq\Big{(}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big{)}I_{d}.$
For the upper bound, it follows from Lemma C.3 and Lemma C.7 that
@@
(C.9) $=\Big{(}\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big{)}I_{d}.$
@@
(C.10) $\nabla b^{*}(t,x)\succeq\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\Big{(}\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big{)}\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}=\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
Combining (C.9) and (C.10) yields (C.8). This competes the proof. ∎
@@
(C.11) $\leq\Big{(}\int_{0}^{1}G\,\mathrm{d}\tau\Big{)}\|x-x^{\prime}\|_{2}=G\|x-x^{\prime}\|_{2},$
@@
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to (C.11). Then applying Gronwall’s inequality [Evans, 2010, Section B.2] completes the proof. ∎
@@
According to Lemma C.6, we find
(C.12) $\|{\mathbb{E}}[X_{1}|X_{t}=x]\|_{2}\leq M_{1}R,\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. For the conditional covariance, using Lemma C.7 implies
(C.13) $\|\operatorname{Cov}(X_{1}|X_{t}=x)\|_{\operatorname{op}}\leq M_{2}\alpha_{t}^{2},\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. In addition, applying Lemma C.8 yields
(C.14) $\|{\mathbb{E}}[X_{1}X_{1}^{T}X_{1}|X_{t}=x]-{\mathbb{E}}[X_{1}^{T}X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}]\|_{2}\leq M_{3}\alpha_{t}^{2}R,$
for each $(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty}$. Substituting (C.12), (C.13) and (C.14) to Lemma C.4 achieves the desired result. ∎
Appendix D Proof of Results in Section 3.3
@@
D.1. Proof of Theorem 3.9
In this section, we prove Theorem 3.9. Specifically, we propose the oracle inequality in Lemma D.1, which decomposes the $L^{2}$-risk into approximation error, the generalization error, and the truncation error. Then we provide an approximation error bound in Lemma D.2. By making a trade-off between three errors, we finally obtain the convergence rate for the velocity estimator, which completes the proof Theorem 3.9.
@@
Lemma D.1 (Oracle inequality for velocity estimation).
@@
(D.1) $\max_{1\leq k\leq d}|b_{k}(t,x)|\leq B_{\operatorname{vel}}R,\quad(t,x)\in[0,T]\times{\mathbb{R}}^{d},$
@@
(D.2) $Y_{t}=b^{*}(t,X_{t})+\varepsilon_{t},\quad X_{t}\sim\mu_{t},\leavevmode\nobreak\ t\sim\operatorname{Unif}[0,T].$
Recall the data set ${\mathcal{S}}=\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$. Then we define the data set corresponding to the regression model (D.2) as $\{(t^{(i)},X_{t}^{(i)},Y_{t}^{(i)})\}$, for which
@@
In this step, we show that the noise term $(\varepsilon_{t}|X_{t}=x)$ in (D.2) is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$, and aim to estimate its variance proxy. According to the definition of stochastic interpolant $X_{t}$ (2.1), we have
@@
which implies from Lemma C.5 that
@@
=$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x,$
where $U\in[0,1]^{d}$ and $\epsilon\sim N(0,I_{d})$ are two independent variables. Then taking expectation on both sides of the equality yields
${\mathbb{E}}[Y_{t}|X_{t}=x]=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U]+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x.$

### v5

@@ D.1. Auxiliary Properties
$=\frac{1}{(2\pi\sigma\alpha_{t})^{d}}\frac{1}{\rho_{t}(x)}\int_{{\mathbb{R}}^{d}}\exp\Big(-\frac{1}{\alpha_{t}^{2}}\|\beta_{t}x_{1}-x\|_{2}^{2}\Big)\exp\Big(-\frac{1}{\sigma^{2}}\|x_{1}-u\|_{2}^{2}\Big)\,\mathrm{d}\nu(u)$
$=\frac{1}{(2\pi)^{d/2}}\Big(\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{d/2}\int_{{\mathbb{R}}^{d}}\exp\Big(-\frac{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\sigma^{2}\alpha_{t}^{2}}\Big\|x_{1}-\frac{\sigma^{2}\beta_{t}x+\alpha_{t}^{2}u}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big\|_{2}^{2}\Big)g(t,x,u)\,\mathrm{d}\nu(u),$
$=N\Big(\frac{\sigma^{2}\beta_{t}x+\alpha_{t}^{2}u}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}},\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)*\nu_{t,x}(u),$
where $g(t,x,u)$ is a function such that $\int\rho_{1|t}(x_{1}|x)\,\mathrm{d}x_{1}=1$ for each $(t,x)\in(0,1)\times{\mathbb{R}}^{d}$, and the measure $\nu_{t,x}$ is defined as $\mathrm{d}\nu_{t,x}(u)=g(t,x,u)\,\mathrm{d}\nu(u)$. It is apparent that $\operatorname{supp}(\nu_{t,x})=\operatorname{supp}(\nu)$. Therefore,
$(X_{1}|X_{t}=x)\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U_{t,x}+\sqrt{\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
where $U_{t,x}\sim\nu_{t,x}$ and $\epsilon\sim N(0,I_{d})$ are two independent random variables. This completes the proof. ∎
Lemma D.6 (Conditional expectation).
@@
Proof of Lemma D.6.
According to Lemma D.5, it holds that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
@@
Lemma D.7 (Conditional covariance).
@@
Proof of Lemma D.7.
It is straightforward from Lemma D.5 that
${\mathbb{E}}[X_{1}|X_{t}=x]=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x,$
@@
$=\Big(\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{2}{\mathbb{E}}[U_{t,x}]{\mathbb{E}}[U_{t,x}]^{T}+\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]x^{T}+\Big(\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{2}xx^{T}.$
On the other hand, using Lemma D.5 deduces
${\mathbb{E}}[X_{1}X_{1}^{T}|X_{t}=x]$ $=\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]x^{T}+\Big(\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{2}xx^{T}$
$\quad+\Big(\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{2}{\mathbb{E}}[U_{t,x}U_{t,x}^{T}]+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
@@
$=\Big(\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big)^{2}\operatorname{Cov}(U_{t,x})+\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
According to Lemma D.5, the random variable $\|U_{t,x}\|_{\infty}\leq 1$ and thus $\operatorname{Cov}(U_{t,x})\preceq dI_{d}$. Consequently,
@@
By a same argument as Lemmas D.6 and D.7, we also have the following inequality. See [Gao et al., 2024, Lemma A.8] for a detailed proof.
Lemma D.8.
@@
D.2. Proof of Results in Section 3.2
@@
where the second equality holds from the definition of stochastic interpolant (2.1), and the last equality is due to Lemma D.2. This completes the proof. ∎
@@
$=\beta_{t}\Big(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big)\Big(\frac{\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]+\frac{\sigma^{2}\beta_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x\Big)+\frac{\dot{\alpha}_{t}}{\alpha_{t}}x$
$=\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]+\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}x$
where the second equality holds from the definition of stochastic interpolant (2.1), and the fourth equality follows from Lemma D.6. This completes the proof. ∎
@@
where the second inequality follows from Lemma D.6, and $M$ is a constant only depending on $d$ and $\sigma$. Applying Gronwall’s inequality [Evans, 2010, Section B.2] yields
@@
First, we observe that
$g^{*}(t,s,x)=x+\int_{t}^{s}b^{*}(\tau,g^{*}(t,\tau,x))\,\mathrm{d}\tau\,.$
Taking the partial derivative with respect to $s$ yields $\partial_{s}g^{*}(t,s,x)=b^{*}(s,g^{*}(t,s,x))$. Since $x\in{\mathbb{B}}_{R}^{\infty}$, by Corollary 3.3 and Proposition 3.2, we get $\|\partial_{s}g^{*}(t,s,x)\|_{2}\leq\sqrt{d}B_{\rm vel}B_{\rm flow}R$.
Now, we consider the partial derivative of $g^{*}(t,s,x)$ with respect to $t$, which leads to
(D.9) $\partial_{t}g^{*}(t,s,x)=-b^{*}(t,x)+\int_{t}^{s}\nabla b^{*}(\tau,g^{*}(t,\tau,x))\partial_{t}g^{*}(t,\tau,x)\,\mathrm{d}\tau\,.$
Note that $\partial_{t}g^{*}(t,t,x):=\partial_{t}g^{*}(t,s,x)|_{s=t}=-b^{*}(t,x)$. Denote by $\zeta(t,s,x):=\partial_{t}g^{*}(t,s,x)$, and further take the partial derivative with respect to $s$ on both sides of (D.9):
$\partial_{s}\zeta(t,s,x)=\nabla b^{*}(s,g^{*}(t,s,x))\zeta(t,s,x)\,.$
Then, it holds that
$\partial_{s}\|\zeta(t,s,x)\|^{2}_{2}=2\big\langle\zeta(t,s,x),\partial_{s}\zeta(t,s,x)\big\rangle=2\big\langle\zeta(t,s,x),\nabla b^{*}(s,g^{*}(t,s,x))\zeta(t,s,x)\big\rangle\,.$
Using Cauchy-Schwarz inequality, we have that
$\partial_{s}\|\zeta(t,s,x)\|^{2}_{2}\leq 2\|\nabla b^{*}(t,x)\|_{\rm op}\|\zeta(t,s,x)\|^{2}_{2}\leq 2G\|\zeta(t,s,x)\|^{2}_{2}\,.$
For any $t\leq s\leq T$, with Gronwall’s inequality [Evans, 2010, Section B.2], it holds that
$\|\zeta(t,s,x)\|^{2}_{2}\leq\|\zeta(t,t,x)\|^{2}_{2}\exp\Big(\int_{t}^{s}2G\,\mathrm{d}\tau\Big)\leq\|b^{*}(x,t)\|_{2}^{2}\exp(2G)\,.$
Thus, we conclude that $\|\partial_{t}g^{*}(t,s,x)\|_{2}\leq\|b^{*}(x,t)\|_{2}\exp(G)\leq\sqrt{d}B_{\rm vel}R\exp(G)$. ∎
@@
(D.10) $\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}\preceq\nabla b^{*}(t,x)\preceq\Big(\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big)I_{d}.$
For the upper bound, it follows from Lemma D.3 and Lemma D.7 that
@@
(D.11) $=\Big(\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}+d\frac{\alpha_{t}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}\Big)I_{d}.$
@@
(D.12) $\nabla b^{*}(t,x)\succeq\frac{\dot{\alpha}_{t}}{\alpha_{t}}I_{d}+\Big(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\Big)\frac{\beta_{t}^{2}}{\alpha_{t}^{2}}\frac{\sigma^{2}\alpha_{t}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}=\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}I_{d}.$
Combining (D.11) and (D.12) yields (D.10). This competes the proof. ∎
@@
(D.13) $\leq\Big(\int_{0}^{1}G\,\mathrm{d}\tau\Big)\|x-x^{\prime}\|_{2}=G\|x-x^{\prime}\|_{2},$
@@
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to (D.13). Then applying Gronwall’s inequality [Evans, 2010, Section B.2] completes the proof. ∎
@@
According to Lemma D.6, we find
(D.14) $\|{\mathbb{E}}[X_{1}|X_{t}=x]\|_{2}\leq M_{1}R,\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. For the conditional covariance, using Lemma D.7 implies
(D.15) $\operatorname{Cov}(X_{1}|X_{t}=x)\|_{\operatorname{op}}\leq M_{2}\alpha_{t}^{2},\quad(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty},$
where $M_{2}$ is a constant only depending on $d$ and $\sigma$. In addition, applying Lemma D.8 yields
(D.16) $\|{\mathbb{E}}[X_{1}X_{1}^{T}X_{1}|X_{t}=x]-{\mathbb{E}}[X_{1}^{T}X_{1}|X_{t}=x]{\mathbb{E}}[X_{1}|X_{t}]\|_{2}\leq M_{3}\alpha_{t}^{2}R,$
for each $(t,x)\in(0,1)\times{\mathbb{B}}_{R}^{\infty}$. Substituting (D.14), (D.15) and (D.16) to Lemma D.4 achieves the desired result. ∎
Appendix E Proof of Results in Section 3.3
@@
E.1. Proof of Theorem 3.10
In this section, we prove Theorem 3.10. Specifically, we propose the oracle inequality in Lemma E.1, which decomposes the $L^{2}$-risk into approximation error, the generalization error, and the truncation error. Then we provide an approximation error bound in Lemma E.2. By making a trade-off between three errors, we finally obtain the convergence rate for the velocity estimator, which completes the proof Theorem 3.10.
@@
Lemma E.1 (Oracle inequality for velocity estimation).
@@
(E.1) $\max_{1\leq k\leq d}|b_{k}(t,x)|\leq B_{\operatorname{vel}}R,\quad(t,x)\in[0,T]\times{\mathbb{R}}^{d},$
@@
(E.2) $Y_{t}=b^{*}(t,X_{t})+\varepsilon_{t},\quad X_{t}\sim\mu_{t},\penalty 10000\ t\sim\operatorname{Unif}[0,T].$
Recall the data set ${\mathcal{S}}=\{(t^{(i)},X_{0}^{(i)},X_{1}^{(i)})\}_{i=1}^{n}$. Then we define the data set corresponding to the regression model (E.2) as $\{(t^{(i)},X_{t}^{(i)},Y_{t}^{(i)})\}$, for which
@@
In this step, we show that the noise term $(\varepsilon_{t}|X_{t}=x)$ in (E.2) is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$, and aim to estimate its variance proxy. According to the definition of stochastic interpolant $X_{t}$ (2.1), we have
@@
which implies from Lemma D.5 that
@@
$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}U_{t,x}+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x,$
where $U_{t,x}\in[0,1]^{d}$ and $\epsilon\sim N(0,I_{d})$ are two independent variables. Then taking expectation on both sides of the equality yields
${\mathbb{E}}[Y_{t}|X_{t}=x]=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x+\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}{\mathbb{E}}[U_{t,x}]+\frac{\sigma^{2}\beta_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})}x.$